# LayerPeeler: Autoregressive Peeling for Layer-wise Image Vectorization

### v1

@@ Abstract.
Image vectorization is a powerful technique that converts raster images into vector graphics, enabling enhanced flexibility and interactivity. However, popular image vectorization tools struggle with occluded regions, producing incomplete or fragmented shapes that hinder editability. While recent advancements have explored rule-based and data-driven layer-wise image vectorization, these methods face limitations in vectorization quality and flexibility. In this paper, we introduce LayerPeeler, a novel layer-wise image vectorization approach that addresses these challenges through a progressive simplification paradigm. The key to LayerPeeler’s success lies in its autoregressive peeling strategy: by identifying and removing the topmost non-occluded layers while recovering underlying content, we generate vector graphics with complete paths and coherent layer structures. Our method leverages vision-language models to construct a layer graph that captures occlusion relationships among elements, enabling precise detection and description for non-occluded layers. These descriptive captions are used as editing instructions for a finetuned image diffusion model to remove the identified layers. To ensure accurate removal, we employ localized attention control that precisely guides the model to target regions while faithfully preserving the surrounding content. To support this, we contribute a large-scale dataset specifically designed for layer peeling tasks. Extensive quantitative and qualitative experiments demonstrate that LayerPeeler significantly outperforms existing techniques, producing vectorization results with superior path semantics, geometric regularity, and visual fidelity. The project page is https://layerpeeler.github.io/.
@@
Figure 1. LayerPeeler enables a layer-wise image vectorization that autoregressively removes topmost layers to recover occluded regions. Each example showcases the key steps in the peeling process, with representative shapes highlighted in the box to demonstrate the structural coherence of the recovered layers.
@@ 1. Introduction
Vector graphics encode visual content through mathematically defined geometric primitives (e.g., paths and Bézier curves), offering resolution-independent rendering and explicit hierarchical representation via layered structures. Unlike raster images, vector representations enable lossless geometric transformations and precise element-wise manipulation, making them widely used in 3D rendering, digital art, and interface design, where structural fidelity and parametric control are essential.
@@
Manually creating complex vector graphics content is a challenging task that demands a steep learning curve and professional software expertise. Image vectorization provides an automated solution to convert raster images into vector formats with minimal manual intervention. However, existing rule-based algorithms (Tian and Günther, 2022) and popular vectorization tools (e.g., Potrace and Adobe Illustrator) suffer from a significant limitation: when processing a partially occluded color region, they often produce incomplete or fragmented vector elements (see Fig. 2). These disjointed components impede intuitive shape manipulation and may introduce visual artifacts when transformed, as the intended semantic structures are not preserved. To address these issues, a layer-aware vectorization method is necessary – one that can reconstruct occluded regions and generate well-structured, editable vector layers while maintaining topological coherence.
One intuitive direction for this is to enhance existing rule-based methods by incorporating awareness of layer occlusion order. For example, Law and Kang (2025) defined energy functions to estimate both pairwise and global occlusion relationships, followed by curvature-based inpainting to reconstruct occluded regions. However, such geometric heuristics often fall short when dealing with layers that contain rich semantic content. An alternative direction lies in data-driven approaches that leverage recent advances in deep learning. Among these, training-free methods (Ma et al., 2022; Hirschorn et al., 2024; Zhou et al., 2024; Wang et al., 2024a) utilize differentiable rasterizers (Li et al., 2020) to render vector shapes into raster images and iteratively optimize shape parameters by minimizing pixel-level losses. They approximate layer appearance either by selecting the largest connected region or by applying SDS loss (Poole et al., 2022) to suppress high-frequency details and obtain layers at different abstraction levels. While capable of achieving high visual fidelity, their strategies for estimating layer order tend to fragment coherent regions or introduce boundary misalignment, requiring substantial post-processing before practical use. Meanwhile, training-based methods (Reddy et al., 2021; Shen and Chen, 2021; Chen et al., 2023; Thamizharasan et al., 2024; Song et al., 2025) aim to learn both layer structure and ordering directly from data via neural networks. However, their performance is constrained by the lack of large-scale, high-quality vector graphics datasets, limiting their ability to generalize across diverse styles and complexities.
@@
Figure 2. Existing algorithmic tools (e.g., Adobe Illustrator) produce vectorizations with incomplete or fragmented shapes. In contrast, our LayerPeeler produces layer-wise vector elements that preserve layer hierarchies, maintain structural integrity, and exhibit clear semantic correspondence.
@@
In this paper, we develop a training-based image vectorization method that converts raster images into vector graphics with strong structural integrity through an iterative, layer-wise approach. The key to the success of LayerPeeler lies in its autoregressive peeling paradigm, which progressively identifies and removes the topmost non-occluded layers while recovering the occluded content beneath – hence the name LayerPeeler. Our method leverages the spatial understanding capabilities of Vision-Language Models (VLMs) to infer occlusion relationships and layer ordering, enabling precise detection and description of non-occluded layers. To mitigate the scarcity of vector graphics data, we harness the strong image priors of image diffusion models. Specifically, we employ a diffusion transformer model fine-tuned with Low-Rank Adaptation (LoRA) (Hu et al., 2021) to interpret layer descriptions and synthesize clean removals of targeted elements. To further enhance visual fidelity and editing precision, we introduce a localized attention control mechanism that constrains modifications to relevant regions during denoising, ensuring seamless integration with the surrounding context for fine-grained element removal.
@@
We demonstrate the effectiveness of our method in the context of Scalable Vector Graphics (SVG), a widely used vector image format. To support the layer peeling task, we construct a new dataset from a large collection of SVG collected from the Internet, enriched with detailed annotations of the topmost layers. This dataset is used to train our image diffusion model for accurate layer-wise modification. We conduct a comprehensive evaluation of LayerPeeler across path semantics, geometric regularity, and visual fidelity. Our experimental results show that LayerPeeler consistently outperforms existing rule-based and data-driven methods both quantitatively and qualitatively. Moreover, we demonstrate the method’s generalization capability across varying SVG styles and complexity levels. Our main contributions are summarized as follows:
@@ 2.1. Image Vectorization
Early work on image vectorization predominantly focused on rule-based approaches, which laid the groundwork for vector graphics representations and creative tools. These methods typically represented vector graphics using meshes (Battiato et al., 2004; Demaret et al., 2006; Swaminarayan and Prasad, 2006; Lecot and Levy, 2006; Liao et al., 2012; Xia et al., 2009; Yang et al., 2015; Sun et al., 2007) or curves (Orzan et al., 2008; Xie et al., 2014; Zhao et al., 2017), offering high-quality vectorization of visual content. However, a key limitation of these algorithmic techniques is their inability to recover occluded content, often resulting in incomplete or fragmented shapes. We refer readers to (Tian and Günther, 2022) for a more comprehensive review of these foundational approaches.
A critical challenge in enabling layer-wise image vectorization is accurately determining the layer order while recovering the occluded content. Rule-based methods, such as (Law and Kang, 2025), approximate ordering by defining energy functions over segmented, non-overlapping regions to construct a global directed graph of pairwise ordering relationships. While computationally efficient, their reliance on geometric inpainting often produces semantically implausible shapes in occluded regions. Recent research has increasingly focused on data-driven approaches that leverage advances in deep learning. Training-free methods (Ma et al., 2022; Hirschorn et al., 2024; Zhou et al., 2024; Wang et al., 2024a) estimate layer appearance either by identifying the largest connected regions or by removing high-frequency details using SDS loss (Poole et al., 2022). These methods achieve high visual fidelity through iterative optimization with DiffVG (Li et al., 2020), but their layer order estimation strategies often fragment connected regions or produce misaligned boundaries that require additional shapes for refinement to improve visual similarity. Training-based methods (Reddy et al., 2021; Shen and Chen, 2021; Chen et al., 2023; Thamizharasan et al., 2024; Song et al., 2025) aim to learn depth priors directly from data, enabling single-pass vectorization through neural networks. However, their effectiveness is constrained by the limited availability of large-scale vector graphics datasets.
@@
Our work focuses on vectorizing non-photorealistic images (Dominici et al., 2020; Du et al., 2023; Favreau et al., 2017; Hoshyari et al., 2018; Yan et al., 2024; Zhang et al., 2009; Su et al., 2021; Kopf and Lischinski, 2011) through an iterative, layer-wise approach. We overcome previous limitations by leveraging VLMs for sophisticated layer-order analysis and by addressing data scarcity through the rich prior knowledge embedded in image diffusion models.
@@ 2.2. Vision-Language Models for Spatial Understanding
Vision-language models have demonstrated remarkable capabilities in understanding and reasoning about visual content, enabling diverse applications across multiple domains. These include visual layout configuration, such as poster design and 3D scene arrangement (Cheng et al., 2024; Yang et al., 2024; Sun et al., 2024), creative content manipulation, like 3D shape editing and motion graphics generation (Ganeshan et al., 2024; Huang et al., 2024; Liu et al., 2024; Ma and Agrawala, 2025), and high-level design understanding and evaluation (Kulits et al., 2024; Wu et al., 2024). Leveraging VLMs’ deep understanding of visual concepts, we employ them to analyze depth and occlusion relationships in input images and construct accurate layer graphs describing the topmost elements. Additionally, through carefully designed visual annotations (Yang et al., 2023; Cai et al., 2024; Lei et al., 2024), we harness their capabilities to annotate layers in vector graphics.
@@ 2.3. Diffusion Models for Image Editing
Image editing with diffusion models can be broadly categorized into training-based and training-free approaches. Training-based methods (Brooks et al., 2023; Wei et al., 2024; Xu et al., 2024a; Yu et al., 2024; Zhang et al., 2023; Zhao et al., 2024) achieve controlled image manipulation by fine-tuning pretrained diffusion models on high-quality datasets, demonstrating impressive editing capabilities through specialized training. Training-free methods offer a more flexible alternative, typically employing a two-stage process of latent inversion followed by denoising (Dong et al., 2023; Hertz et al., 2022; Meng et al., 2021; Tumanyan et al., 2023). Recent work has focused on improving the accuracy of the inversion stage through advanced sampling techniques (Avrahami et al., 2024; Cao et al., 2023; Li et al., 2023; Tewel et al., 2024; Xu et al., 2024b). Notably, attention modulation (Ju et al., 2023; Lin et al., 2024; Miyake et al., 2025; Mokady et al., 2023; Wang et al., 2024b) has emerged as an effective tool in training-free methods due to its simplicity and precise control over editing operations. Building on these advances, our approach combines both paradigms – we first train a specialized model using a curated layer removal dataset, then manipulate the attention masks during inference to achieve accurate layer deletion.
Figure 3. System diagram of LayerPeeler. Given a rasterized source image (parrot), our system leverages vision-language models to analyze the image across three dimensions: (1) constructing a layer graph that captures occlusion relationships among color regions, (2) identifying the topmost layers and generating a global prompt that describes them, and (3) decomposing the topmost layers into instance-level elements, each with an associated prompt and bounding box. The prompts and source image are encoded through text and image encoders to obtain $c_{e}$ and $x_{\mathrm{src}}$, respectively, while a noise vector $z_{T}$ is initialized. These inputs are fed into our LoRA-finetuned diffusion model, which synthesizes an image with the specified layers removed. The bounding boxes guide attention masks within joint attention modules. This layer removal process operates autoregressively, iteratively identifying and removing topmost layers until a complete layer-wise decomposition is achieved.
@@ 3. LayerPeeler
In this section, we first introduce the definition and construction of a layer graph using VLMs to accurately detect non-occluded shapes (Section 3.1). We then describe the detailed architecture of our image diffusion model along with the localized attention sculpting technique for precise layer removal (Section 3.2). Next, we outline our dataset construction process, which enables diffusion models to perform the layer peeling task effectively (Section 3.3). Finally, we present the autoregressive inference strategy and vectorization method (Section 3.4).
@@ 3.2. Layer Peeling with Diffusion Models
As no existing pretrained image generation model supports the fine-grained image editing required for our task, i.e., the precise removal of individual layers, we fine-tune an image diffusion model to enable text-guided layer peeling. We adopt the state-of-the-art Diffusion Transformer (DiT) architecture (Peebles and Xie, 2023), specifically FLUX.1-dev, as our base model. This choice is motivated by its superior generation quality compared to previous models and its robust text-conditioning capabilities, enabled by its integration with the T5 text encoder (Raffel et al., 2020).
During training, we sample triplets $(\mathcal{P}_{e},I_{\mathrm{src}},I_{\mathrm{tar}})$ from our custom-built dataset (Section 3.3). The editing prompt $\mathcal{P}_{e}$, source image $I_{\mathrm{src}}$, and target image $I_{\mathrm{tar}}$ are transformed into tokens using text and image encoders, yielding $c_{e}$, $x_{\mathrm{src}}$, and $x_{\mathrm{tar}}$, respectively. For a given timestep $t\in[0,1]$, we sample a noise vector $\epsilon\sim\mathcal{N}(0,I)$ and compute the noised latent representation as $z=(1-t)x_{\mathrm{tar}}+t\epsilon$. The noisy latent $z$ is then concatenated with the source token sequence $x_{\mathrm{src}}$ along the sequence dimension. To encourage the model to learn spatial correspondences between the source and target images, we apply the same positional encoding to both $z$ and $x_{\mathrm{tar}}$, following the approach of Huang et al. (2025). We adopt the Conditional Flow Matching (CFM) loss from SD3 (Esser et al., 2024) to optimize the model during training:
@@
where $u_{t}(z|\epsilon)$ denotes the conditional vector field, and $v_{\Theta}(z,x_{\mathrm{src}},c_{e},t)$ represents the velocity field parameterized by the neural network’s weights $\Theta$. To avoid the computational cost of full-parameter finetuning, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2021), a parameter-efficient training strategy. LoRA freezes the pretrained model weights and injects trainable low-rank update matrices: $\Delta\mathbf{W}=\mathbf{A}\mathbf{B}$ into specified layers, where $\mathbf{A}\in\mathbb{R}^{m\times r}$, $\mathbf{B}\in\mathbb{R}^{r\times n}$, and the rank $r\ll\min(m,n)$. The final weight matrices are computed as $\mathbf{W}^{\prime}=\mathbf{W}+\Delta\mathbf{W}$.

### v2

@@ Abstract.
Image vectorization is a powerful technique that converts raster images into vector graphics, enabling enhanced flexibility and interactivity. However, popular image vectorization tools struggle with occluded regions, producing incomplete or fragmented shapes that hinder editability. While recent advancements have explored optimization-based and learning-based layer-wise image vectorization, these methods face limitations in vectorization quality and flexibility. In this paper, we introduce LayerPeeler, a novel layer-wise image vectorization approach that addresses these challenges through a progressive simplification paradigm. The key to LayerPeeler’s success lies in its autoregressive peeling strategy: by identifying and removing the topmost non-occluded layers while recovering underlying content, we generate vector graphics with complete paths and coherent layer structures. Our method leverages vision-language models to construct a layer graph that captures occlusion relationships among elements, enabling precise detection and description for non-occluded layers. These descriptive captions are used as editing instructions for a finetuned image diffusion model to remove the identified layers. To ensure accurate removal, we employ localized attention control that precisely guides the model to target regions while faithfully preserving the surrounding content. To support this, we contribute a large-scale dataset specifically designed for layer peeling tasks. Extensive quantitative and qualitative experiments demonstrate that LayerPeeler significantly outperforms existing techniques, producing vectorization results with superior path semantics, geometric regularity, and visual fidelity. Our code and dataset will be available at https://layerpeeler.github.io/.
@@
Figure 1. LayerPeeler enables a layer-wise image vectorization that autoregressively removes topmost layers to recover occluded regions. Each example showcases the key steps in the peeling process, with representative shapes highlighted in the box to demonstrate the structural coherence of the recovered layers. The icon images are from ©SVGRepo.
@@ 1. Introduction
Vector graphics encode visual content through mathematically defined geometric primitives, offering resolution-independent rendering and explicit hierarchical representation via layered structures. Unlike raster images, vector representations enable lossless geometric transformations and precise element-wise manipulation, making them widely used in 3D rendering, digital art, and interface design, where structural fidelity and parametric control are essential.
@@
Manually creating complex vector graphics content is a challenging task that demands a steep learning curve and professional software expertise. Image vectorization provides an automated solution to convert raster images into vector formats with minimal manual intervention. However, existing rule-based algorithms (tian2022survey) and popular vectorization tools (e.g., Potrace and Adobe Illustrator) suffer from a significant limitation: when processing a partially occluded color region, they often produce incomplete or fragmented vector elements (see Fig. 2). These disjointed components impede intuitive shape manipulation and may introduce visual artifacts when transformed, as the intended semantic structures are not preserved. To address these issues, a layer-aware vectorization method is necessary – one that can reconstruct occluded regions and generate well-structured, editable vector layers while maintaining topological coherence.
To this end, one line of work pursues optimization-based geometric methods (favreau2017photo2clipart; du2023image; law2025image). Starting from a segmentation map, they optimize an objective that trades off image fidelity against structural regularization (e.g., layer count and smooth boundaries) and infer layer order from perceptual cues (e.g., T-/X-junctions) or energy-based criteria. However, with shape completion based on geometric features instead of semantics, reconstructions are frequently oversimplified and sometimes fail entirely. Beyond purely geometric formulations, a second optimization-based line leverages differentiable rendering to optimize vector primitives (ma2022towards; hirschorn2024optimize; zhou2024segmentation; wang2024layered). These methods use a differentiable rasterizer (Li:2020:DVG) to minimize reconstruction or perceptual losses, and synthesize layers jointly or progressively from background (deepest) to foreground, which can yield detailed reconstructions but often over-partition shapes and produce messy layer structures. In parallel, learning-based methods (reddy2021im2vec; shen2021clipgen; chen2023editable; thamizharasan2024vecfusion; song2025layertracer) aim to predict layered vectors in a single pass, but are bottlenecked by the scarcity of large-scale vector datasets, limiting generalization across styles and complexities.
@@
Figure 2. Existing vectorization tools (e.g., Adobe Illustrator) produce vectorizations with incomplete or fragmented shapes. In contrast, our LayerPeeler produces layer-wise vector elements that preserve layer hierarchies, maintain structural integrity, and exhibit clear semantic correspondence. The icon image is from ©SVGRepo.
@@
In this paper, we develop a learning-based image vectorization method that converts raster images into vector graphics with strong structural integrity through an iterative, layer-wise approach. The key to the success of LayerPeeler lies in its autoregressive peeling paradigm, which progressively identifies and removes the topmost non-occluded layers while recovering the occluded content beneath – hence the name LayerPeeler. Our method leverages the spatial understanding capabilities of Vision-Language Models (VLMs) to infer occlusion relationships and layer ordering, enabling precise detection and description of non-occluded layers. To mitigate the scarcity of vector graphics data, we harness the strong image priors of image diffusion models. Specifically, we employ a diffusion transformer model fine-tuned with Low-Rank Adaptation (LoRA) (hu2021lora) to interpret layer descriptions and synthesize clean removals of targeted elements. To further enhance visual fidelity and editing precision, we introduce a localized attention control mechanism that constrains modifications to relevant regions during denoising, ensuring seamless integration with the surrounding context for fine-grained element removal.
@@
We demonstrate the effectiveness of our method in the context of Scalable Vector Graphics (SVG), a widely used vector image format. To support the layer peeling task, we construct a new dataset from a large collection of SVG collected from the Internet, enriched with detailed annotations of the topmost layers. This dataset is used to train our image diffusion model for accurate layer-wise modification. We conduct a comprehensive evaluation of LayerPeeler across path semantics, geometric regularity, and visual fidelity. Our experimental results show that LayerPeeler consistently outperforms existing optimization-based and learning-based methods both quantitatively and qualitatively. Moreover, we demonstrate the method’s generalization capability across varying SVG styles and complexity levels. Our main contributions are summarized as follows:
@@ 2.1. Image Vectorization
Early work on image vectorization predominantly focused on rule-based approaches, which laid the groundwork for vector graphics representations and creative tools. These methods typically represented vector graphics using meshes (battiato2004svg; demaret2006image; swaminarayan2006rapid; lecot2006ardeco; liao2012subdivision; xia2009patch; yang2015effective; sun2007image) or curves (orzan2008diffusion; xie2014hierarchical; zhao2017inverse), offering high-quality vectorization of visual content. However, a key limitation of these algorithmic techniques is their inability to recover occluded content, often resulting in incomplete or fragmented shapes. We refer readers to (tian2022survey) for a more comprehensive review of these foundational approaches.
A critical challenge in enabling layer-wise image vectorization is accurately determining the layer order while recovering the occluded content. To this end, optimization-based methods cast vectorization as energy minimization that trades off reconstruction fidelity against structural regularity. Within this family, geometric optimization approaches follow a common recipe: starting from a segmentation map, they formulate an objective that explicitly balances a reconstruction term against structural regularizers (e.g., penalizing layer count, encouraging semi-transparency and convex/smooth boundaries), infer layer order from perceptual cues (e.g., T-/X-junctions) or energy-based criteria, and complete occlusions via geometric or variational models, yielding a small number of editable layers (favreau2017photo2clipart; entem2018structuring; du2023image; law2025image). However, because completion relies on geometric features rather than semantics, these methods can oversimplify occluded regions, struggle with non-convex objects, and are sensitive to segmentation quality, leading to semantically implausible layers or orderings in complex scenes. In parallel, differentiable rendering optimization methods iteratively refine vector graphics by minimizing reconstruction or perceptual losses through a differentiable rasterizer (Li:2020:DVG), achieving high visual fidelity but frequently at the cost of an excessive number of layers and fragmented or misaligned regions (ma2022towards; hirschorn2024optimize; zhou2024segmentation; wang2024layered).
More recently, learning-based methods leverage deep networks to directly map raster images to vector representations. By learning depth priors and structural patterns from data, these models enable single-pass vectorization (reddy2021im2vec; shen2021clipgen; chen2023editable; thamizharasan2024vecfusion; song2025layertracer). However, their performance remains constrained by the limited availability of large-scale vector graphics datasets, which hinders generalization across diverse styles and domains.
@@
Our work focuses on vectorizing non-photorealistic images (dominici2020polyfit; du2023image; favreau2017photo2clipart; hoshyari2018perception; yan2024deep; Zhang2009VectorizingCA; su2021marvel; kopf2011depixelizing), which play an important role in design applications, through an iterative, layer-wise approach. We overcome previous limitations by leveraging VLMs for sophisticated layer-order analysis and by addressing data scarcity through the rich prior knowledge embedded in image diffusion models.
@@ 2.2. Vision-Language Models for Spatial Understanding
Vision-language models have demonstrated remarkable capabilities in understanding and reasoning about visual content, enabling diverse applications across multiple domains. These include visual layout configuration, such as poster design and 3D scene arrangement (cheng2024graphic; yang2024posterllava; sun2024layoutvlm), creative content manipulation, like 3D shape editing and motion graphics generation (ganeshan2024parsel; huang2024blenderalchemy; liu2024logomotion; ma2025mover), and high-level design understanding and evaluation (kulits2024re; wu2024gpt). Leveraging VLMs’ deep understanding of visual concepts, we employ them to analyze depth and occlusion relationships in input images and construct accurate layer graphs describing the topmost elements. Additionally, through carefully designed visual annotations (yang2023set; cai2024vip; lei2024scaffolding), we harness their capabilities to annotate layers in vector graphics.
@@ 2.3. Diffusion Models for Image Editing
Image editing with diffusion models can be broadly categorized into learning-based and training-free approaches. learning-based methods (brooks2023instructpix2pix; wei2024omniedit; xu2024insightedit; yu2024anyedit; zhang2023magicbrush; zhao2024ultraedit) achieve controlled image manipulation by fine-tuning pretrained diffusion models on high-quality datasets, demonstrating impressive editing capabilities through specialized training. Training-free methods offer a more flexible alternative, typically employing a two-stage process of latent inversion followed by denoising (dong2023prompt; hertz2022prompt; meng2021sdedit; tumanyan2023plug). Recent work has focused on improving the accuracy of the inversion stage through advanced sampling techniques (avrahami2024stable; cao2023masactrl; li2023stylediffusion; tewel2024add; xu2024headrouter). Notably, attention modulation (ju2023direct; lin2024schedule; miyake2025negative; mokady2023null; wang2024taming) has emerged as an effective tool in training-free methods due to its simplicity and precise control over editing operations. Building on these advances, our approach combines both paradigms: we first train a specialized model using a curated layer removal dataset, then manipulate the attention masks during inference to achieve accurate layer deletion.
Figure 3. System diagram of LayerPeeler. Given a rasterized source image (parrot), our system leverages vision-language models to analyze the image across three dimensions: (1) constructing a layer graph that captures occlusion relationships among color regions, (2) identifying the topmost layers and generating a global prompt $\mathcal{P}_{g}$ that describes them (then forming the editing prompt $\mathcal{P}_{e}$ by prepending the word “remove” to $\mathcal{P}_{g}$), and (3) decomposing the topmost layers into instance-level elements, each with an associated prompt and bounding box. The prompts and source image are encoded through text and image encoders to obtain $c_{e}$ and $x_{\mathrm{src}}$, respectively, while a noise vector $z_{T}$ is initialized. These inputs are fed into our LoRA-finetuned diffusion model, which synthesizes an image with the specified layers removed. The bounding boxes guide attention masks within joint attention modules. This layer removal process operates autoregressively, iteratively identifying and removing topmost layers until a complete layer-wise decomposition is achieved. The icon image is from ©SVGRepo.
@@ 3. LayerPeeler
In this section, we first introduce the definition and construction of a layer graph using VLMs to accurately detect non-occluded shapes (Section 3.1). We then describe the detailed architecture of our image diffusion model along with the localized attention sculpting technique for precise layer removal (Section 3.2). Next, we present the autoregressive inference strategy and vectorization method (Section 3.3). Finally, we outline our dataset construction process, which enables diffusion models to perform the layer peeling task effectively (Section 3.4).
@@ 3.2. Layer Peeling with Diffusion Models
As no existing pretrained image generation model supports the fine-grained image editing required for our task, i.e., the precise removal of individual layers, we fine-tune an image diffusion model to enable text-guided layer peeling. We adopt the state-of-the-art Diffusion Transformer (DiT) architecture (peebles2023scalable), specifically FLUX.1-dev, as our base model. This choice is motivated by its superior generation quality compared to previous models and its robust text-conditioning capabilities, enabled by its integration with the T5 text encoder (raffel2020exploring).
During training, we sample triplets $(\mathcal{P}_{e},I_{\mathrm{src}},I_{\mathrm{tar}})$ from our custom-built dataset (Section 3.4). The editing prompt $\mathcal{P}_{e}$, source image $I_{\mathrm{src}}$, and target image $I_{\mathrm{tar}}$ are transformed into tokens using text and image encoders, yielding $c_{e}$, $x_{\mathrm{src}}$, and $x_{\mathrm{tar}}$, respectively. For a given timestep $t\in[0,1]$, we sample a noise vector $\epsilon\sim\mathcal{N}(0,I)$ and compute the noised latent representation as $z=(1-t)x_{\mathrm{tar}}+t\epsilon$. The noisy latent $z$ is then concatenated with the source token sequence $x_{\mathrm{src}}$ along the sequence dimension. To encourage the model to learn spatial correspondences between the source and target images, we apply the same positional encoding to both $z$ and $x_{\mathrm{tar}}$, following the approach of huang2025photodoodle. We adopt the Conditional Flow Matching (CFM) loss from SD3 (esser2024scaling) to optimize the model during training:
@@
where $u_{t}(z|\epsilon)$ denotes the conditional vector field, and $v_{\Theta}(z,x_{\mathrm{src}},c_{e},t)$ represents the velocity field parameterized by the neural network’s weights $\Theta$. To avoid the computational cost of full-parameter finetuning, we employ Low-Rank Adaptation (LoRA) (hu2021lora), a parameter-efficient training strategy. LoRA freezes the pretrained model weights and injects trainable low-rank update matrices: $\Delta\mathbf{W}=\mathbf{A}\mathbf{B}$ into specified layers, where $\mathbf{A}\in\mathbb{R}^{m\times r}$, $\mathbf{B}\in\mathbb{R}^{r\times n}$, and the rank $r\ll\min(m,n)$. The final weight matrices are computed as $\mathbf{W}^{\prime}=\mathbf{W}+\Delta\mathbf{W}$.
@@
3.3. Inference
Inference begins by sending the source image to VLMs to get the tuple $\{\mathcal{G},\mathcal{P}_{g},\{\mathcal{B}_{i},\mathcal{P}_{i}\}_{i=1}^{N}\}$ and initializing a random noise vector $z_{T}$, which serves as the starting point for the denoising process. The latents of the source image, denoted as $x_{\mathrm{src}}$, are kept noise-free during denoising steps to provide spatial correspondence and preserve high-frequency details. Once the diffusion model completes the layer-peeling operation, it outputs an image in which the specified topmost layer has been removed. To isolate the removed region, we compute a pixel-level difference map between the source and output images, apply thresholding ($\rho=20$ by default) to create a binary mask, and use this mask to extract the removed region from the source image. The extracted region is then vectorized using an off-the-shelf vectorization tool. This entire pipeline is executed autoregressively: after each layer is removed and vectorized, the resulting output image becomes the new source image for the next iteration. The process continues until the image becomes completely white. Finally, we stack all the vectorized layers in reverse order to reconstruct the complete SVG.
Figure 4. The visual annotation used for non-occluded layer captioning. The icon image is from ©SVGRepo.